# Inflationary Dynamics in Einstein-Gauss-Bonnet Gravity Using New Slow-Roll Approximations considering Generalised Reheating

### v1

@@ VII Reheating Parameters
After the end of the inflationary epoch, the universe goes into a super cold mode. In a standard cold inflationary scenario, the only degree of freedom that can survive after the inflation is the inflation field itself. However to repopulate the universe an era of reheating is required[80, 81, 82, 82], during which the inflaton field transfers its energy to other degrees of freedom, reviving the universe from a super-cooled state to a hot thermal bath of relativistic particles. The concept of reheating was first proposed in [81]. In the standard cold inflationary scenario, various methods of reheating have been suggested in the literature. One such method is perturbative decay, where the inflaton field reaches the bottom of its potential and begins to decay into other elementary particles[83, 84, 85, 86]. The resulting particles interact and reach equilibrium at a temperature known as the reheating temperature ($T_{re}$). More sophisticated methods include parametric resonance, which is a non-perturbative method, tachyonic instability[87], and preheating[88, 89]. The initial stage of reheating is often attributed to the preheating phase, which is more efficient than perturbative reheating because it involves exponential decay, generating a large number of particles.
@@
Assuming, $w_{re}$ to be constant during the reheating epoch, the energy density of the universe at the end of inflation can be expressed in terms of scale factor through $\rho\propto a^{-3(1+w)}$, formulated as follows:
@@
The subscript $end$ signify the end of inflation, while $re$ stands for the end of the reheating period. Using the end of inflation condition($\epsilon=1$), we find $\rho_{end}=(3/2)V_{end}$ and substituting it in eq. 7.60, we find
$N_{re}=\frac{1}{3(1+w_{re})}\ln\left(\frac{\rho_{end}}{\rho_{re}}\right)=\frac{1}{3(1+w_{re})}\ln\left(\frac{3}{2}\frac{V_{end}}{\rho_{re}}\right),$ (7.61)
Also we know:
@@
Where, $g_{re}$ denotes the number of relativistic species at the end of reheating.
Using (7.61) and (7.62) and following [90, 92, 93], we can express $T_{re}$ and $N_{re}$ as :
$N_{re}=\frac{1}{3(1+w_{re})}\ln\left(\frac{30\cdot\frac{3}{2}V_{end}}{\pi^{2}g_{re}T_{re}^{4}}\right)$ (7.63)
@@
$T_{re}=\left[\left(\frac{43}{11g_{re}}\right)^{\frac{1}{3}}\frac{a_{0}T_{0}}{k}H_{k}e^{-N_{k}}\left[\frac{3^{2}\cdot 5V_{end}}{\pi^{2}g_{re}}\right]^{-\frac{1}{3(1+w_{re})}}\right]^{\frac{3(1+w_{re})}{3w_{re}-1}}.$ (7.67)
From 7.66 and 7.67 it is evident that to evaluate $N_{re}$ and $T_{re}$ first we require to calculate the $H_{k}$, $N_{k}$ (e-folds during the inflation) and $V_{end}$ (value of potential at the end of inflation). Utilizing the tensor-to-scalar ratio with scalar power spectrum we can write
@@
From Eq.(7.68), it’s clear that $H_{k}$ is a function of the tensor to scalar ratio ($r$). As we know from the CMB observation, tensor to scalar ratio has only upper bound. To precisely analyze the dynamics of reheating we need to express $H_{k}$ in terms of the spectral index ($n_{s}$). This can be done from the relation between $r,n_{s}$ and $N_{k}$[91, 78]. As mentioned in [90] varying $A_{s}$ with $n_{s}$ has negligible effect on the reheating. we keep $A_{s}(k_{0})=2.0989\times 10^{-9}$.
Figure 7.4: Plots illustrating $N_{re}$ and $T_{re}$ for different choices of $\omega_{re}$ are presented. Left Panel: The results for slow-roll approximation I is represented. Right Panel: The results for slow-roll approximation II is represented. The light yellow shaded area is for the $2-\sigma$ bounds, while the blue shaded area represents the $1-\sigma$ bounds on $n_{s}$ from Planck’18 [13]. Finally, the self-explanatory color coding for $w_{re}$ is given in the inset of the plots.
The variation of $N_{re}$ and $T_{re}$ for the PLP model with $n=2$ and $q=1$ for various values of $w_{re}$ is shown in Fig. 7.4. The convergence points on the $T_{re}$ and $N_{re}$ plots represent instantaneous reheating, with $N_{re}=0$. One can of course do this exercise for all the 16 combinations for this class of inflationary model. But the essence of the findings can be encapsulated in one of these figures such as fig. 7.4.
@@ VIII Conclusions and Discussions
With future observations like CMB-S4 [94] and COrE [95] with promising prospects to measure the spectral tilt very precisely ($\Delta n_{s}\sim 0.002$), and with future possibilities to constrain the primordial tensor modes, a revisit to theoretically well-motivated models are in order. Not only that, the conceptual development in the modified gravity sector in the last few years has opened a window to study several interesting features to study. In the case of EGB gravity, new developments such as the new slow roll approximations methods as proposed in [1], beg a more detailed analysis. In this work, we had a threefold objective:
  • Rescue a theoretically well-motivated model using a non-standard gravitational background of EGB gravity which we show can be done successfully.
@@
Thus, using this prescription, we have summarised a path to study inflationary dynamics in the EGB gravity. We have shown the class of PLP inflationary dynamics in the EGB background can successfully give observationally astute results even keeping the generalized reheating in mind.
A few more studies that we plan to conduct in the EGB inflationary the framework will be to see how the dynamics of Warm inflation get incorporated [96, 97]. Production of Primordial Black Holes and Gravitational Waves in case of warm inflationary dynamics [98, 99, 100, 101] in the EGB background, utilizing this new prescription could be a really interesting path to follow. The authors would like to come to these questions in future studies.
@@ Acknowledgments
The authors would like to thank Mostafizur Rahman, Arun Kumar, and Abolhassan Mohammadi for useful discussion. The work of MRG is supported by the Department of Science and Technology(DST), Government of India under the Grant Agreement number IF18-PH-228 (INSPIRE Faculty Award) and by the Science and Engineering Research Board(SERB), DST, Government of India under the Grant Agreement number CRG/2020/004347(CoreResearch Grant). MRG would like to thank IUCAA, Pune, India, for their hospitality during the visit as an associate, when the project was initiated. MRG would also like to thank OMEG Institute, Soongsil University, Republic of Korea, for their hospitality, where the paper’s final part was completed.
@@ A.1 New Slow Roll Approximation I
×$\times\left(\xi_{1}^{2}\xi_{2}^{2}\text{sech}^{4}\left(\xi_{2}\phi\right)\left(\frac{\phi^{n}}{m^{n}+\phi^{n}}\right)^{2q}-\frac{1}{2}3n\xi_{1}\xi_{2}qm^{n}\phi^{-n-1}\text{sech}^{2}\left(\xi_{2}\phi\right)\left(\frac{\phi^{n}}{m^{n}+\phi^{n}}\right)^{q+1}+\frac{9}{8}\right)$
@@
×$\times\bigg{(}81nq\phi m^{n}\left(m^{n}+\phi^{n}\right)\left(m^{n}(nq-1)-(n+1)\phi^{n}\right)+512m^{n}nq\phi\left(\frac{\phi^{n}}{m^{n}+\phi^{n}}\right)^{6q}(m^{n}+\phi^{n})\left(m^{n}(nq-1)-(n+1)\phi^{n}\right)$
@@ A.2 New Slow Roll Approximation II
×$\times\xi_{1}\xi_{2}\left(2\xi_{2}\phi\left(m^{n}+\phi^{n}\right)\tanh\left(\xi_{2}\phi\right)-nqm^{n}\right)\Bigg{)}\Bigg{)}\Bigg{/}27\Bigg{(}\phi\left(m^{n}+\phi^{n}\right)-2n\xi_{1}\xi_{2}qm^{n}\text{sech}^{2}\left(\xi_{2}\phi\right)\left(\frac{\phi^{n}}{m^{n}+\phi^{n}}\right)^{q}\Bigg{)}{}^{3}$
@@
×$\quad\times\left(-2\xi_{2}\phi\left(m^{n}+\phi^{n}\right)\tanh\left(\xi_{2}\phi\right)+m^{n}(3nq-1)-(n+1)\phi^{n}\right)-32n\xi_{1}^{3}\xi_{2}^{3}q\phi m^{n}\left(m^{n}+\phi^{n}\right)\text{sech}^{6}\left(\xi_{2}\phi\right)\left(\frac{\phi^{n}}{m^{n}+\phi^{n}}\right)^{3q}\times$

### v3

@@ VII Reheating Parameters
After the end of the inflationary epoch, the universe goes into a super-cold mode. In a standard cold inflationary scenario, the only degree of freedom that can survive after the inflation is the inflation field itself. However, to repopulate the universe, an era of reheating is required[93, 94, 95, 95, 96, 97, 98], during which the inflaton field transfers its energy to other degrees of freedom, reviving the universe from a super-cooled state to a hot thermal bath of relativistic particles. The concept of reheating was first proposed in [94]. In the standard cold inflationary scenario, various methods of reheating have been suggested in the literature. One such method is perturbative decay, where the inflaton field reaches the bottom of its potential and begins to decay into other elementary particles[99, 100, 101, 102]. The resulting particles interact and reach equilibrium at a temperature known as the reheating temperature ($T_{re}$). More sophisticated methods include parametric resonance, which is a non-perturbative method, tachyonic instability[103], and preheating[104, 105]. The initial stage of reheating is often attributed to the preheating phase, which is more efficient than perturbative reheating because it involves exponential decay, generating a large number of particles.
@@
Assuming $w_{re}$ to be constant during the reheating epoch, the energy density of the universe at the end of inflation can be expressed in terms of the scale factor through $\rho\propto a^{-3(1+w)}$, formulated as follows:
@@
The subscript endend signifies the end of inflation, while $re$ stands for the end of the reheating period. Using the end of inflation condition($\varepsilon=1$).
Here crucial point is to calculate the $\rho_{end}$ for the EGB gravity. In standard cold inflationary scenarios, it is straightforward to calculate $\rho_{end}$ by employing $\omega=\frac{P}{\rho}=-\frac{1}{3}$. However, in EGB gravity, the end of inflation might not occur at $\omega=-\frac{1}{3}$, and it can be slightly different. From Eq. 2.3 (keeping $U_{0}=1/2$)
$3H^{2}=\frac{\psi^{2}}{2}+V+12\xi_{,\phi}\psi H^{3}$ (7.40)
Comparing this with the standard cold inflationary scenario where $3H^{2}=\rho_{\phi}=\frac{\psi^{2}}{2}+V$ we can compute the extra contribution to the energy density due to the modified gravity($\rho_{egb}=\xi_{,\phi}\psi H^{3}$).
$3H^{2}=\rho_{\phi}+\rho_{egb}$ (7.41)
Similarly from Eq. 2.4 we can write
$2\dot{H}=-\left(\psi^{2}-4H^{2}\xi_{,\phi\phi}\psi^{2}-4H^{2}\xi_{,\phi}\dot{\psi}+4H^{3}\xi_{,\phi}\psi-8\xi_{,\phi}\psi H\dot{H}\right)$ (7.42)
Now comparing this with $2\dot{H}=-(P_{\phi}+\rho_{\phi})$ we can compute the extra contribution to the pressure due to the presence of the EGB gravity. We write
$2\dot{H}=-\Big{(}\left(P_{\phi}+P_{egb}\right)+\left(\rho_{\phi}+\rho_{egb}\right)\big{)}$ (7.43)
where, $P_{egb}=-4H^{2}\xi,_{\phi\phi}\psi^{2}-4H^{3}\xi,_{\phi}\dot{\psi}-\frac{8}{3}H^{3}\xi,_{\phi}\psi-8\xi,_{\phi}\psi H\dot{H}$, now we calculate the $\rho_{\phi}$ in terms of the inflationary potential from $\omega_{\phi}=\frac{P_{\phi}}{\rho_{\phi}}$ at the end of inflation.
$\rho_{\phi}=V\left(\frac{2}{1-\omega_{\phi}}\right)$ (7.44)
Also, we are writing $\frac{\rho_{egb}}{\rho_{\phi}}=\alpha$, this allows us to write the total energy density at the end of inflation as:
$\rho_{\rm total}=\rho_{\phi}+\rho_{egb}=\rho_{\phi}(1+\alpha)=\beta V,~~~~~~~{\rm where}~~~~\beta=\left(\frac{2}{1-\omega_{\phi}}\right)\left(1+\alpha\right)$ (7.45)
Substituting it in eq. (7.39), we find
$N_{re}=\frac{1}{3(1+w_{re})}\ln\left(\frac{\rho_{end}}{\rho_{re}}\right)=\frac{1}{3(1+w_{re})}\ln\left({\color[rgb]{0,0,1}\beta}\frac{V_{end}}{\rho_{re}}\right),$ (7.46)
Also, we know:
@@
Where $g_{re}$ denotes the number of relativistic species at the end of reheating. Using (7.46) and (7.47) and following [106, 108, 109], we can express $T_{re}$ and $N_{re}$ as :
$N_{re}=\frac{1}{3(1+w_{re})}\ln\left(\frac{30~{\beta}~V_{end}}{\pi^{2}g_{re}T_{re}^{4}}\right)$ (7.48)
@@
$T_{re}=\left[\left(\frac{43}{11g_{re}}\right)^{\frac{1}{3}}\frac{a_{0}T_{0}}{k}H_{k}e^{-N_{k}}\left[\frac{30~{\beta}~V_{end}}{\pi^{2}g_{re}}\right]^{-\frac{1}{3(1+w_{re})}}\right]^{\frac{3(1+w_{re})}{3w_{re}-1}}.$ (7.52)
It is noteworthy to mention that the expressions of $N_{re}$ and $T_{re}$ obtained here are general expressions. Which are valid for the EGB gravity, also all the information of EGB gravity is encoded in the $\beta$, $n_{s},N_{k}$, and $H_{k}$ through the coupling $\xi(\phi)$. Evolution and the converging point of $N_{re}$ and $T_{re}$ are different for both the slow roll approximation and the exact solution. This is because the evolution of $n_{s},N_{k}$, and $H_{k}$ is slightly different in approximations and numerical solutions. However, the order of the quantity $\beta$ is always around $\mathcal{O}(1)$, which was also the case in the standard cold inflation scenario. Thus, the inclusion of $\beta$ has a negligible effect on the behavior of $N_{re}$ and $T_{re}$. Also, the first term in the square bracket in Eq. 7.51 has a contribution from $\beta$, but this contribution is also minimal. From Eq.(7.51) and (7.52) it is evident that to evaluate $N_{re}$ and $T_{re}$, first, we require to calculate the $H_{k}$, $N_{k}$ (e-folds during the inflation) and $V_{end}$ (value of the potential at the end of inflation). Utilizing the tensor-to-scalar ratio with the scalar power spectrum,m we can write
@@
From Eq.(7.53), it’s clear that $H_{k}$ is a function of the tensor-to-scalar ratio ($r$). As we know from the CMB observation, tensor to scalar ratio has only an upper bound. To precisely analyze the dynamics of reheating, we need to express $H_{k}$ in terms of the spectral index ($n_{s}$). This can be done from the relation between $r,n_{s}$ and $N_{k}$[107, 90]. As mentioned in [106], varying $A_{s}$ with $n_{s}$ has negligible effect on the reheating, so we keep $A_{s}(k_{0})=2.0989\times 10^{-9}$.
Figure 7.6: Plots illustrating $N_{re}$ and $T_{re}$ for different choices of $\omega_{re}$ are presented. Left Panel: The results for slow-roll approximation I are represented. Right Panel: The results for slow-roll approximation II are represented. The light blue shaded area is for the $2-\sigma$ bounds, while the light green shaded area represents the $1-\sigma$ bounds on $n_{s}$ from Planck’18 [14]. Finally, the self-explanatory color coding for $w_{re}$ is given in the inset of the plots.
Figure 7.7: Plots illustrating $N_{re}$ and $T_{re}$ for different choices of $\omega_{re}$ by solving system ( 2.6) numerically for $n=2$ and $q=1$. Plot legends are self-explanatory.
The variation of $N_{re}$ and $T_{re}$ for the PLP model with $n=2$ and $q=1$ for various values of $w_{re}$ is shown in Fig. (7.6). In Fig. 7.7 we present the reheating analysis for the exact solution obtained by numerically solving Eq. ( 2.6). The convergence points on the $T_{re}$ and $N_{re}$ plots represent instantaneous reheating, with $N_{re}=0$. One can, of course, do this exercise for all 16 combinations for this class of inflationary model. But the essence of the findings can be encapsulated in one of these figures, such as fig. (7.6) and (7.7).
@@ VIII Conclusions and Discussions
With future observations like CMB-S4 [110] and COrE [111], which have promising prospects to measure the spectral tilt very precisely ($\Delta n_{s}\sim 0.002$), and with future possibilities to constrain the primordial tensor modes, a revisit to theoretically well-motivated models is in order. Not only that, the conceptual development in the modified gravity sector in the last few years has opened a window to study several interesting features. In the case of EGB gravity, new developments such as the new slow roll approximation methods proposed in [1], beg for a more detailed analysis. In this work, we had a threefold objective:
  • Rescue a theoretically well-motivated model using a non-standard gravitational background of EGB gravity, which we show can be done successfully.
@@
Thus, using this prescription, we have summarised a path to study inflationary dynamics in the EGB gravity. We have shown that the class of PLP inflationary dynamics in the EGB background can successfully yield observationally astute results, even keeping the generalized reheating in mind. In our analysis, we found field values at the start of inflation in Slow Roll Approximation-I are close to the field values obtained by numerically solving Eq. 2.6 (see Table 1.2,1.2 and 1.3). Thus, we conclude that in our study, Slow Roll Approximation-I is more accurate than the Slow Roll Approximation-II
A few more studies that we plan to conduct in the EGB inflationary framework will be to see how the dynamics of Warm inflation get incorporated [112, 113]. Production of Primordial Black Holes and Gravitational Waves in case of warm inflationary dynamics [114, 115, 116, 117] in the EGB background, utilizing this new prescription could be a really interesting path to follow. The authors would like to address these questions in future studies.
@@ Acknowledgments
The authors would like to thank Mostafizur Rahman, Arun Kumar, and Abolhassan Mohammadi for useful discussions. The work of MRG is supported by the Department of Science and Technology(DST), Government of India under the Grant Agreement number IF18-PH-228 (INSPIRE Faculty Award) and by the Science and Engineering Research Board(SERB), DST, Government of India, under the Grant Agreement number CRG/2022/004120(CoreResearch Grant). MRG would like to thank IUCAA, Pune, India, for their hospitality during the visit as an associate when the project was initiated. MRG would also like to thank OMEG Institute, Soongsil University, Republic of Korea, for their hospitality, where the paper’s final part was completed.
@@ A.1 New Slow Roll Approximation I
$\times\left(\xi_{1}^{2}\xi_{2}^{2}\text{sech}^{4}\left(\xi_{2}\phi\right)\left(\frac{\phi^{n}}{m^{n}+\phi^{n}}\right)^{2q}-\frac{1}{2}3n\xi_{1}\xi_{2}qm^{n}\phi^{-n-1}\text{sech}^{2}\left(\xi_{2}\phi\right)\left(\frac{\phi^{n}}{m^{n}+\phi^{n}}\right)^{q+1}+\frac{9}{8}\right)$
@@
$\times\bigg{(}81nq\phi m^{n}\left(m^{n}+\phi^{n}\right)\left(m^{n}(nq-1)-(n+1)\phi^{n}\right)+512m^{n}nq\phi\left(\frac{\phi^{n}}{m^{n}+\phi^{n}}\right)^{6q}(m^{n}+\phi^{n})\left(m^{n}(nq-1)-(n+1)\phi^{n}\right)$
@@ A.2 New Slow Roll Approximation II
$\times\xi_{1}\xi_{2}\left(2\xi_{2}\phi\left(m^{n}+\phi^{n}\right)\tanh\left(\xi_{2}\phi\right)-nqm^{n}\right)\Bigg{)}\Bigg{)}\Bigg{/}27\Bigg{(}\phi\left(m^{n}+\phi^{n}\right)-2n\xi_{1}\xi_{2}qm^{n}\text{sech}^{2}\left(\xi_{2}\phi\right)\left(\frac{\phi^{n}}{m^{n}+\phi^{n}}\right)^{q}\Bigg{)}{}^{3}$
@@
$\quad\times\left(-2\xi_{2}\phi\left(m^{n}+\phi^{n}\right)\tanh\left(\xi_{2}\phi\right)+m^{n}(3nq-1)-(n+1)\phi^{n}\right)-32n\xi_{1}^{3}\xi_{2}^{3}q\phi m^{n}\left(m^{n}+\phi^{n}\right)\text{sech}^{6}\left(\xi_{2}\phi\right)\left(\frac{\phi^{n}}{m^{n}+\phi^{n}}\right)^{3q}\times$
@@
A.3 Field values in Slow Roll approximations and Numerical analysis
Table 1.1: Table for field value at start of inflation ($\phi_{\rm i}$) and end of inflation ($\phi_{\rm e}$) for different values of potential parameters when $\Delta N=60$ in Slow-Roll I
$q$ $n$ $\phi_{\rm i}$ $\phi_{\rm e}$
1 8.8622 0.8987
1 2 8.6941 1.2516
3 8.6167 1.3598
4 8.5966 1.3836
1 9.1927 1.4516
2 2 8.8128 1.7501
3 8.6536 1.7205
4 8.6131 1.6510
1 9.5287 1.7497
3 2 8.9218 2.0874
3 8.6833 1.9572
4 8.6243 1.8212
1 9.8713 1.8422
4 2 9.0266 2.3325
3 8.7102 2.1378
4 8.6333 1.9498
Table 1.2: Table for field value at start of inflation ($\phi_{\rm i}$) and end of inflation ($\phi_{\rm e}$) for different values of potential parameters when $\Delta N=60$ in Slow-Roll II
$q$ $n$ $\phi_{\rm i}$ $\phi_{\rm e}$
1 8.9850 3.3626
1 2 8.7967 3.5634
3 8.7181 3.5727
4 8.7009 3.5579
1 9.2961 3.1765
2 2 8.8963 3.5800
3 8.7391 3.6010
4 8.7048 3.5744
1 9.6231 2.9812
3 2 8.9955 3.5919
3 8.7601 3.6261
4 8.7088 3.5897
1 9.9600 2.7781
4 2 9.0942 3.6001
3 8.7309 3.6487
4 8.7127 3.6041
Table 1.3: Table for field value at start of inflation ($\phi_{\rm i}$) and end of inflation ($\phi_{\rm e}$) for different values of potential parameters when $\Delta N=60$ obtained by numerically solving Eq. 2.6.
$q$ $n$ $\phi_{\rm i}$ $\phi_{\rm e}$
1 8.9022 0.4532
1 2 8.6732 0.8638
3 8.6013 1.0831
4 8.5851 1.1783
1 9.1311 0.6928
2 2 8.7753 1.1894
3 8.6273 1.3588
4 8.5934 1.4015
1 9.4492 0.8829
3 2 8.8875 1.3964
3 8.6513 1.5272
4 8.6000 1.5341
1 9.7838 1.0458
4 2 8.9744 1.5573
3 8.6743 1.6533
4 8.6060 1.6304